# EnKF-C user guide version 2.32.0

### v11

@@ 2.13 Use of innovation statistics for model validation
EnKF-C can calculate innovation statistics for validating a model against observations only, without data assimilation. The pre-requisites are (i) observations and (ii) model dump readable by the code, and possibly (iii) auxiliary files for projecting the model state to observation space (e.g. grid specs and mean SSH). To get the innovation statistics one needs to:
@@ 2.14 Bias correction
[…]
@@ 2.16.2 Memory footprint
Object Typical size1 PREP CALC UPDATE
observation array 6 GB
super-observation array2 1.5 GB
ensemble observations2 $\mathcal{H}(\mathbf{E})$ 5 GB $\times$ 2
single grid2,3 0.4 GB
observation K-D trees2 0.85 GB
one 3D model field 1 GB
ensemble of one horizontal field 2 GB
transform array (I/O, EnKF only) 22 GB4
@@
The memory footprint of PREP is defined by the size of the (original, before superobing) observation structure array and, in some cases, by the size of curvilinear grids. The memory usage by curvilinear grids is much reduced by parsing them into K-D trees (default from v1.101.4; the only option since v1.106.0) rather than into binary trees. Because PREP is not parallelised (mainly due to lack of robust and efficient parallel analogue of qsort procedure), in practice its memory footprint is rarely a problem.
@@
For the EnKF, the footprint of UPDATE is mainly defined by the size of the array of simultaneously updated horizontal model fields. The number of simultaneously updated fields is defined by parameter FIELDBUFFERSIZE. Note that larger values of FIELDBUFFERSIZE increase computational effectiveness by reducing the number of reads of and interpolations within $\mathbf{X}_{5}$ arrays. For the EnOI, the footprint of UPDATE is insensitive to FIELDBUFFERSIZE (which should be set to 1), and is defined by the size of the ensemble of horizontal model fields.
@@ 2.17 Possible problems / FAQ
    In the innovation stats report at the tail of CALC log I get some entries “N/A” for instrument tags.
    These entries show stats for super-observations obtained by collating observations from different instruments. To avoid collating such observations run PREP with option
    --no-superobing-across-instruments. This became default from v1.115.0.
  3. 3.
@@
  4. 4.
    CALC starts calculating transforms swiftly, but then slows down almost to a standstill.
    This can happen with large systems when the MPI communication becomes clogged. Recompile CALC with flag -DTW_VIAFILE to communicate transforms via filesystem.
  5. 5.
@@
  6. 6.
    In the main cycles of CALC and UPDATE some CPUs finish quickly (display 100 %), while other continue for much longer.
    The jobs look unbalanced, try compiling with flag -DUSE_MPIQUEUE.
  7. 7.

### v12

@@ 2.13 Use of innovation statistics for model validation
EnKF-C can calculate innovation statistics for validating a model against observations only, without data assimilation. The prerequisites are (i) observations and (ii) model dump readable by the code, and possibly (iii) auxiliary files for projecting the model state to observation space (e.g. grid specs and mean SSH). To get the innovation statistics one needs to:
@@ 2.14 Bias correction
where $\lambda$ is the forgetting factor, $0<\lambda<1,\;1-\lambda\ll 1$, and $\boldsymbol{\sigma}\sim\sigma_{0}N(0,1)$, where $\sigma_{0}$ is the error standard deviation of $\mathbf{x}$. This can be specified for a model variable via entry RANDOMISE in the corresponding section of the model parameter file (sec. 2.4.2).
@@ 2.16.2 Memory footprint
Object Typical size1 Typical size2 PREP CALC UPDATE
observation array 6 GB 8 GB
super-observation array2 1.5 GB 3 GB
ensemble observations3 $\mathcal{H}(\mathbf{E})$ 5 GB $\times$ 2 21 GB $\times$ 2
single grid3 0.4 GB4 ($\times$ 2) 1.6 GB ($\times$ 3)
observation K-D trees3 0.85 GB 2 GB
one 3D model field 1 GB 4.6 GB
ensemble of one horizontal field 2 GB 10 GB
transform array (file) 22 GB5 ($\times$ 2) 29 GB5 ($\times$ 3)
@@
(2) for EnKF/GOSI9 system ($4322\times 3606\times 75+5$ grid, 24 dynamic + 144 static members, $8\cdot 10^{7}$ observations, $3.2\cdot 10^{7}$ super-observations)
@@
The memory footprint of PREP is defined by the size of the (original, before superobing) observation structure array and, in some cases, by the size of curvilinear grids. The memory usage by curvilinear grids is much reduced by parsing them into K-D trees (default from v1.101.4; the only option since v1.106.0) rather than into binary trees. Because PREP is not parallelised (mainly due to lack of robust and efficient parallel analogue of qsort procedure), in practice its memory footprint is rarely a problem (when running from a master script).
@@
For the EnKF, the footprint of UPDATE is mainly defined by the size of the array of horizontal model fields times the number of simultaneously updated fields. The number of simultaneously updated fields is defined by parameter FIELDBUFFERSIZE. Note that larger values of FIELDBUFFERSIZE increase computational effectiveness by reducing the number of reads of and interpolations within $\mathbf{X}_{5}$ arrays. For the EnOI, the footprint of UPDATE is insensitive to FIELDBUFFERSIZE (which should be set to 1), and is defined by the size of the ensemble of horizontal model fields.
For very large models like 1/12-degree NEMO in GOSI9 configuration the ensemble of horizontal fields ($\sim 10$ GB) can exceed the average memory per core (“slot”). This requires reducing the number of used slots per NUMA node when running UPDATE.
@@ 2.17 Possible problems / FAQ
  3. 3.
@@
  4. 4.
@@
Chapter 3 Setting up a DA system
This chapter lists the main steps for setting up a geophysical ensemble DA (forecasting) system. Specifically, it concerns (but not limited to) ocean DA systems.
A DA system is built based on a geophysical model. Normally, a running model is sufficient for setting up a DA system with EnKF-C.
3.1 Initial considerations
Typical ocean DA systems aim at constraining mesoscale circulation because the available amount of observations and quality of ocean models make it impossible to reliably constrain smaller scales. Considering that (1) the characteristic error doubling time at mesoscale is of order of a month; (2) unlike variational DA, there is no need for large observation windows in ensemble DA; but (3) smaller cycle lengths incur additional computational overheads, the reasonable DA cycle lengths in ocean DA systems are somewhere between a day and a week.
3.2 EnKF systems
  1. 1.
    Decide on the ensemble size.
    In our experience, best performance is achieved with hybrid systems, which also allow smaller ensembles than pure EnKF systems. For a pure EnKF system, a typical recommended ensemble size is about 100, although often it can be reduced without substantial deterioration in performance.
  2. 2.
    Create initial ensemble.
    For the EnKF one needs a “wide” initial ensemble such that, ideally, the true system state at the initial time might be considered as a sample from it. The detailed composition of the initial EnKF ensemble is not important because it will be “forgotten” after the spin-up.
    In the context of ocean forecasting the initial ensemble can be created by conducting a long model run (hindcast) and saving model restarts around the time of the year (season) of the initial time. The length of this hindcast can be from about 5 to 30 years or more, typically about 10 years. The number of model restarts (ensemble members) saved each year is determined by achieving the desired ensemble size. Note that if the system is supposed to run with randomly perturbed forcing then it is possible to duplicate some members as they quickly diverge after starting the system.
    A challenging part of the initial hindcast is to obtain possibly larger subsurface variability. Typically, the subsurface states of ocean models are substantially different from the true state; therefore larger subsurface variability is important for quicker reduction of initial biases (which can take years). Larger subsurface variability can be obtained by longer initial hundcasts; by conducting several hindcasts initialised from different ocean state products; and by perturbing mixing and advection settings.
  3. 3.
    Decide on the Mean Dynamic Topography (MDT) product.
    For global systems it normally should be one of the available products that is or can be used in calculating sea level anomaly (SLA) observations. For systems with local area models (LAMs) it should be the product used in the system providing the open boundary conditions.
    In some cases the model can be substantially biased about MDT estimates. Thes biases project on subsurface state and can cause significant drifts and biases. In these cases it may be necessary to use the native MDT calculated as the average sea level during the initial hindcast.
  4. 4.
    Choose observation time binning, localisation radii and R-factors.
    Generally, it is preferable to use time-average model outputs for calculating forecast observations. The time intervals chosen for particular model variables should be based on the their temporal variability. Ocean circulation at mesoscale evolves very slowly. Therefore, the timing for assimilating SLA, temperature and salinity (T and S) observations is not really critical. It seems natural to calculate forecast observations using daily average model outputs but it is also possible to use longer intervals or assimilate synchronously using model restarts.
    In contrast, SST exhibits a strong diurnal (daily) cycle. Therefore, SST needs to be matched against model outputs on a sub-daily scale. Further on, because it is impossible today to reliably model the daytime SST, one needs (1) to restrict assimilation of SST observations to nighttime only and (2) to ensure that no daytime model SST is contributing to the forecast observations. Once resticted to nighttime, the SST observations will be matched in DA with SST from the closest in time model dump. It is reasonable to use 1- or 2-hour time intervals for SST model outputs.
    Localisation radii and R-factors should be chosen to provide robust inversions in local analyses and sensible increments. The same concerns the magnitude of K-factor (sec. 2.7.3).
  5. 5.
    Choose moderation settings.
    The values of multiplicative inflation and magnitude of RPS (secs. 2.8.1,2.7.4) are set to ensure robust evolution of the DAS, with as small adjustments to the KF (no inflation, no relaxation) as possible. To minimise long-term damage from multiplicative inflation it is important to keep it minimal and cap for each element of the state vector by the magnitude of spread reduction (sec. 2.8.1). (Capped inflation is a default setting in EnKF-C.) See section 3.4.1 for an example of setings from an operational global ocean forecasting system.
  6. 6.
    DA cycle design.
    This involves positioning of observation window relative to analysis. For EnKF systems it is rather straightforward. Assuming that the system is optimal enough, asynchronous DA should work well for an asymmetric cycle (when the analysis is not centred about the observation window). In practice, centred observation window may yield somewhat better performance but it is unlikely to be justified by 1.5x increase of the cost of model propagation, which typically constitutes 90–95 % of the total cost of the EnKF. Therefore, it is natural to set observation window to the interval from previous to current analysis.
  7. 7.
    Spin-up of the system
    The system starts from an unconstrained state; therefore, the initial corrections are expected to be rather unbalanced. A practical way to spin-up the system is to increase the supposedly optimal R-factors by a large multiple and gradually reduce it in the first several cycles. For example, one can use the multiples $2^{7-i}$ for cycle $i,i=0,\dots,6$ for the first 7 cycles.
  8. 8.
    Correct the initial state
    The initial state is represented by the ensemble average of the initial ensemble generated from a long run of the free model. This initial state can be substantially biased about SLA, T and S observations. In this case it is a good idea to correct it (i.e. subtract the global mean bias from ensemble fields for T and S and add to MDT for SSH); otherwhile it can take a long time (of order of a year) to get it reduced by DA.
3.3 EnOI
  1. 1.
    Create static ensemble.
    Unlike the inital ensemble for an EnKF system, which gets “forgotten” over time, the static ensemble in an EnOI system does not evolve and therefore continuosly affects the performance of the system. Because of its static nature, it cannot represent the “errors of the day” and should aim at representing large scale variability.
    The EnOI ensemble is created based on a long model run. In contrast to the EnKF, the main outputs should be not model restarts but time average fields for the updated variables. The ensemble should not contain long term trends including seasonal trends. This can be accomplished by generating ensemble members as the difference between time averaged model fields over two time windows of different width centered about each other. The particular lengths of these windows can be optimised depending on the model resolution and domain. Typically, for ocean forecasting, the inner window can be from 1-2 days to a week, and the outer window from 1-2 weeks to 1-3 months. In practice, it is convenient to save 1-day model average fields during the hindcast and then generate the ensemble based on this output.
    The size of the static ensemble can be typically about 200. Larger ensembles start affecting the cost of matrix inversions (that scale as the size of the matrix cubed) while bringing diminishing benefits for the system performance.
    Note that for ocean EnKF systems it is desirable to exclude the atmospheric component of SSH (“inverse barometer”) from the saved fields; otherwise the ensemble SSH anomalies would contain this extra variability, which negatively affects the balances between model fields in the ensemble. (Although the suggested way of calculating ensemble anomalies as the difference of time average in two centred windows can largely eliminate the atmospheric variability in SSH.)
  2. 2.
    DA cycle design.
    The EnOI is essentially an interpolation technique and, as such, produces increments for the centre of the observation window. Further, because EnOI systems propagate only one instance of the model, the cost of model propagation is not as critical as for the EnKF. It is therefore natural to centre the observation window about the analysis time.
    With FGAT approach this means that at each DA cycle the model needs to be propagated for the half of DA cycle to the start of the next observation window and then further for the full cycle to the end of the observaiton window, i.e. for 1.5 DA cycle in total. The model needs to save restart at the time of the next analysis, i.e. after propagating for 1.0 DA cycle.
  3. 3.
    DA settings.
    Generally, EnOI systems reach best performance with somewhat stronger fit to observations than EnKF systems (with smaller R-factors and larger localisation radii than in EnKF systems with ensembles of similar spread). Note that the ensemble spread depends on the ensemble generation technique and settings, which involves additional tuning by means of a multiple to R-factors.
3.4 Hybrid EnKF/EnOI
Today hybrid EnKF/EnOI systems represent the state-of-the art in the ensemble based ocean forecasting. For example, OceanMAPS v4.0 and v4.1 are global (75 S to 75 N) eddy resolving (0.1) ocean forecasting systems run by Australian Bureau of Meteorology. They are hybrid systems with 48 dynamic and 144 static ensemble members. These systems considerably outperform otherwise similar EnKF systems with 96 ensemble members at half of their computational cost.
Setting up a hybrid system involves the following steps.
  1. 1.
    Generate static and dynamic ensemble.
    Follow steps described in the previous sections for generating EnKF and EnOI ensembles.
  2. 2.
    Decide on the scaling coefficient GAMMA of the static ensemble.
    In practice OceanMAPS does not show much sensitivity to GAMMA.
  3. 3.
    DA cycle design.
    It seems natural to follow the design described for the EnKF. In situations when computing expenses are not an issue it is also possible to use observation window symmetric about analysis.
3.4.1 Settings in OceanMAPS v4.1
Ensemble size: 48 dynamic members, 144 static members
Cycle length: 1 d
Observation window: [1,0][-1,0] d about analysis
Observation binning: SLA 1 d, SST 2 h, T 1 d, S 1 d
R-factors: SLA 4.5, SST 48, T 18, S 18
Loc. support radii (km): SLA 175, SST 150, T 450, S 450
ALPHA: 0.70
GAMMA: 0.25
Inflation: 1.02 1
K-factor: 2